OpenStudy (anonymous):

Can someone help with this problem. It is uber aggravating nd I just want help already. Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (reemii):

3sqrt(2x) + xsqrt(8x) - 5sqrt(18x) = 3sqrt(2x) + 2xsqrt(2x) - 15sqrt(2x) = -10 sqrt(2x) ??? Nakim !! He added 3 to 2x to -15 and obtained -10x. He created a few "x" out of nothing. Correct: sqrt(2x) (3 + 2x - 15) = sqrt(2x) (-12 + 2x).

OpenStudy (anonymous):

Omg your amazing:)